# Quantum many-body analysis of a spin-2 bosons with two-body inelastic decay

### v1

@@ IV.1 Without quadratic Zeeman effect
Figure 1(a) shows the time evolution of the average particle number $\langle\hat{N}\rangle$ and the normalized total magnetization $F^{2}_{\textrm{norm}}$. The average particle number rapidly decreases in the early stages, since the initial state contains an abundance of states with $F<2N$ subject to inelastic decay. After several hundred milliseconds, the states with $F<2N$ are depleted and the decrease in $\langle\hat{N}\rangle$ becomes slow; $\langle\hat{N}\rangle$ finally converges to $\simeq 3$ over long periods. The normalized total magnetization $F^{2}_{\textrm{norm}}$ increases monotonically, as shown in Fig. 1(b), and finally converges to unity, which indicates that the final state is the fully magnetized steady state, given by Eq. (18) with Eq. (17). Because of the rotational symmetry of the system, the expectation values $\langle\hat{F}_{x}\rangle,\langle\hat{F}_{y}\rangle$, and $\langle\hat{F}_{z}\rangle$ of the magnetization remain zero as the system evolves.
@@
Figures 1(b) and 1(c) show the distributions of the number of particles $P_{N}$ and the magnetization $Q_{m}^{(N)}$ at different times. In Fig. 1(b), it can be seen that the distribution becomes broad and its peak shifts toward smaller particle numbers due to inelastic loss. The distribution at $t=1.76~\textrm{s}$ in Fig. 1(c) is similar to the final distribution for the steady state, which has a peak at $N=2$. Figure 1(c) displays the distribution $Q_{m}^{(N)}$ for $N=6$ and $N=16$. We note that $Q_{m}^{(N)}$ is broadened by inelastic loss of particles with $m_{1}+m_{2}\neq 0$, where $m_{1}$ and $m_{2}$ are the magnetic quantum numbers of the lost particles. As a result, a statistical mixture of states with different $m$ is generated.
For comparison, the dotted lines in Fig. 1(c) show
$Q_{m}=\frac{1}{2^{4N}}\binom{4N}{2N+m}$, (38)
which is the distribution of the mean-field states
$|\Psi_{\textrm{MF}}^{(N)}(\phi)\rangle=\hat{R}(\phi)|F=2N,F_{z}=2N\rangle,$ (39)
where $\hat{R}(\phi)=e^{-i\phi\hat{F}_{x}}e^{-i\pi/2\hat{F}_{y}}$ is the rotation operator. These mean-field states correspond to the fully polarized states in the $x$-$y$ directions. The dotted lines for $N=6$ agree well with the results of the numerical simulation, which indicates that the steady state can be approximated by a mixture of $\hat{\rho}_{N}$ as
$\hat{\rho}_{N}=\frac{1}{2\pi}\int d\phi|\Psi_{\textrm{MF}}^{(N)}(\phi)\rangle\langle\Psi_{\textrm{MF}}^{(N)}(\phi)|.$ (40)
This state is a statistical mixture of mean-field states polarized in the $x$-$y$ directions. On the other hand, the distribution $Q_{m}^{(16)}$ in Fig. 1(c) deviates substantially from the dotted line, i.e., the state $\hat{\rho}_{16}$ differs from the mean-field-like state in Eq. (40). However, the probability of observing this state is very small; $P_{16}\sim 10^{-5}$ (see (iii) in Fig. 1(b)).
@@
Figure 2: Time evolution for $q/\hbar=90.6~\textrm{Hz}$ with and without quenching of $q$. In the case with quenching, $q$ is suddenly changed to 0 at $t=14.3~\textrm{ms}$. (a) Transverse magnetization $S_{\perp}$ with quenching (solid line) and without quenching (dotted line). (b) Average particle number $\langle\hat{N}\rangle$ (blue or dark-gray lines) and normalized total magnetization $F^{2}_{\textrm{norm}}$ (red or light-gray lines) with quenching (solid lines) and without quenching (dotted lines). (c) Particle number distribution at $t=1.76~\textrm{s}$ with quench (filled boxes) and without quench (striped boxes). (d) Distribution $Q_{m}^{(N)}$ of $z$-magnetization for $N=6$ and $N=16$ at $t=1.76\textrm{s}$. The dotted lines show Eq. (38).
@@ IV.2 With quadratic Zeeman effect
We apply a quadratic Zeeman term of $q/\hbar=90.6~\textrm{Hz}$ to the system, which corresponds to a magnetic field of $0.178~\textrm{G}$. The dotted line in Fig. 2(a) shows the transverse magnetization $S_{\perp}$. In the early stages of evolution, $S_{\perp}$ rapidly becomes large because of the quadratic Zeeman effect, followed by oscillation. Such behavior has also been observed experimentally [21]. Subsequently, $S_{\perp}$ decays to a small value due to the particle loss. The dotted lines in Fig. 2(b) show the time evolution of the average particle number $\langle\hat{N}\rangle$ and the total magnetization $F^{2}_{\textrm{norm}}$. The average particle number monotonically decreases with time and converges to a small value, since the steady state consists of only a limited set of states, as given by Eqs. (32) and (33). The normalized total magnetization $F^{2}_{\textrm{norm}}$ first oscillates and then increases toward unity. Although the final steady state should have $F^{2}_{\textrm{norm}}=1$, the convergence is very slow, as shown in Fig. 2(b). The striped boxes in Fig. 2(c) show the particle number distribution $P_{N}$ at $t=1.76~\textrm{s}$, which has a peak at $N=2$. This distribution is similar to that for $q=0$ in Fig. 1(b).
Next, we consider the case in which the quadratic Zeeman coefficient is suddenly quenched to $q=0$ at $t=14.3~\textrm{ms}$, at which $S_{\perp}$ reaches its first peak. The solid line in Fig. 2(a) shows that $S_{\perp}$ remains large for a long time. The solid lines in Fig. 2(b) show that $\langle\hat{N}\rangle$ also remains large ($\simeq 13$) and $F^{2}_{\textrm{norm}}$ rapidly converges to unity. The particle number distribution has a peak at $N=16$, as shown by the filled boxes in Fig. 2(c). These behaviors contrast sharply with those in the case without quenching. This is because the fully polarized states in Eq. (17) generated by the quadratic Zeeman term remain entirely in the steady state after $q$ is quenched.
Figure 2(d) shows the distribution $Q_{m}^{(N)}$ of the $z$-magnetization at $t=1.76~\textrm{s}$, which is almost the steady state. For $N=6$, the distribution is in agreement with that for the mean-field-like state in Eq. (40), shown by the dotted line. However, for $N=16$, for which the probability $P_{N}$ is largest, the distribution is nontrivial, exhibiting multiple peaks at $m=0$ and $m=\pm 2$. This characteristic distribution, in which even $m$ is enhanced and odd $m$ is suppressed, suggests a quantum mechanical superposition of fully polarized states oriented in opposite directions ($\theta=-\pi/2$ and $\pi/2$) on the $x$-$y$ plane. This is understood from the fact that the Schödinger-cat-like state given by
$\hat{R}(\pi/2)+\hat{R}(-\pi/2)](\hat{a}_{2}^{\dagger})^{N}|vac\rangle$ (41)
has a distribution $Q_{m}^{(N)}$ that is zero for odd $m$ and is proportional to Eq. (38) for even $m$ [28, 29]. Interestingly, such a nontrivial quantum superposition is achieved despite the system being an open system subject to particle dissipation.
@@ V Conclusions
We have investigated the dynamics of spin-2 bosons with atomic dissipation. In the previous study [21], it was shown experimentally and numerically that the dissipation can enhance the transverse magnetization in the time evolution, and the results of the mean-field calculations were found to agree well with the experiments. By contrast, the present work focused on the quantum many-body nature of a system with a small number of particles. We proved that the steady states for this system are restricted to statistical mixtures of states with maximum total spins, given by Eq. (18) with Eq. (17). We performed numerical simulations using the master equation with and without the quadratic Zeeman term. Without the quadratic Zeeman term, we found that the system indeed reached a steady state with the maximum total spins. When the quadratic Zeeman term was removed in the early stages of time evolution, a relatively large number of particles were found to remain in the steady state. We concluded that this state resembles a Schrödinger-cat-like state, formed by the superposition of mean-field-like states. This suggests that particle dissipation not only destroys the quantum properties but also creates nontrivial quantum many-body states. Although in the present study we considered $^{87}\textrm{Rb}$ atoms, other atomic species, such as $^{23}\textrm{Na}$ atoms, may be experimentally suitable, since their inelastic decay rates are much larger [30, 31, 32] and steady states can be reached much faster.
ACKNOWLEDGMENTS

### v2

@@ IV.1 Without quadratic Zeeman effect
First, we numerically solve the master equation for $q=0$. Figure 1(a) shows the time evolution of the average particle number $\langle\hat{N}\rangle$ and the normalized total magnetization $F^{2}_{\mathrm{norm}}$. The average particle number rapidly decreases in the early stage, since the initial density operator contains an abundance of states with $F<2N$, which are subject to inelastic decay. After several hundred milliseconds, the states with $F<2N$ are depleted and the decrease in $\langle\hat{N}\rangle$ becomes slow; $\langle\hat{N}\rangle$ finally converges to $\simeq 3$ over long periods. The normalized total magnetization $F^{2}_{\mathrm{norm}}$ increases monotonically, as shown in Fig. 1(a), and finally converges to unity; this behavior indicates that the final state is the fully magnetized steady state, given by Eq. (26). Because of the rotational symmetry of the system, the expectation values $\langle\hat{F}_{x}\rangle,\langle\hat{F}_{y}\rangle$, and $\langle\hat{F}_{z}\rangle$ of the magnetization remain zero as the system evolves.
The decrease in $\langle\hat{N}\rangle$ in Fig. 1(a) can roughly be estimated as follows. Assuming that $\mathrm{Tr}\left[\hat{A}_{\mathcal{FM}}^{\dagger}\hat{A}_{\mathcal{FM}}\hat{\rho}\right$] in Eq. (34) is of the order of $\langle\hat{N}\rangle^{2}$, we have $d\langle\hat{N}\rangle/dt\sim-b\langle\hat{N}\rangle^{2}/V^{\mathrm{eff}}$, where $b$ is a parameter comparable to the values of $b_{0}$ and $b_{2}$. This differential equation is solved to give
$\langle\hat{N}\rangle=\frac{1}{\frac{b}{V^{\mathrm{eff}}}t+\langle\hat{N}\rangle_{0}^{-1}},$ (48)
where $\langle\hat{N}\rangle_{0}$ is the initial average particle number. Fitting the numerical results for $\langle\hat{N}\rangle$ to Eq. (48) for $0\leq t\leq 0.22~\mathrm{s}$ with $\langle\hat{N}\rangle_{0}=N_{\mathrm{ini}}$, we find the value of $b$ to be $b\simeq 0.0081\times 10^{-11}\mathrm{cm}^{3}/\mathrm{s}$. The inset in Fig. 1(a) compares Eq. (48) with the numerically obtained average particle number $\langle\hat{N}\rangle$ on a log-log scale, where $T=bt/V^{\mathrm{eff}}+N_{\mathrm{ini}}^{-1}$. Equation (48) well describes the initial decrease in $\langle\hat{N}\rangle$. The numerical result deviates from Eq. (48) for large $T$, since the fully magnetized atoms $\simeq 3$ remain in the steady state, which is not included in Eq. (48).
Figure 1(b) displays the particle number distributions $P_{N}$ at different times. As time progresses, we observe that the distribution becomes broad, and that its peak moves toward smaller particle numbers because of inelastic loss. For large $t$, the particle number distribution converges to a shape resembling a Poisson distribution restricted to even integers,
$P_{N}=\frac{\eta^{N}}{N!\cosh(\eta)},$ (49)
with a peak at $N=2$. Fitting the observed distribution to a Poisson distribution yields $\eta\simeq 3.08$, which is shown as the dotted line in Fig. 1(b). The mechanism that the observed particle number distribution resembles a Poisson distribution merits further study.
@@
Since the initial state in Eq. (42) is an eigenstate of $\hat{F}_{z}$ with an eigenvalue 0, the initial magnetization distribution $Q_{m}$ in the subspace of $N=N_{\mathrm{ini}}=20$ is localized at $m=0$ (i.e., $Q_{m}=\delta_{m,0}$). At later times, $Q_{m}$ is broadened by inelastic loss of particles with $m_{1}+m_{2}\neq 0$, where $m_{1}$ and $m_{2}$ are the magnetic quantum numbers for the lost particles. As a result, a statistical mixture of states with different $m$ is generated in the subspaces of $N<N_{\mathrm{ini}}$. Figure 1(c) shows $Q_{m}$ at $t=8.8$ s, at which the distributions have almost converged to those of the final steady state. The distribution $Q_{m}$ in the $N=N_{\mathrm{ini}}=20$ subspace is still localized at $m=0$, and is broader distributions in subspaces with smaller $N$.
For comparison, we consider the $N$-particle fully polarized mean-field state,
$|\Psi_{\mathrm{MF}}^{(N)}(\varphi)\rangle=\hat{R}(\varphi)|F=2N,F_{z}=2N\rangle,$ (50)
where $\hat{R}(\varphi)=e^{-i\varphi\hat{F}_{x}}e^{-i\pi/2\hat{F}_{y}}$ is the rotation operator. This state is fully polarized in the $x$-$y$ direction with an azimuthal angle $\varphi$, and all $N$ atoms occupy the same spin state with a distribution
$Q^{\mathrm{MF}}_{m}=\frac{1}{2^{4N}}\binom{4N}{2N+m}$. (51)
The density operator after the time evolution is a statistical mixture of different $m$ (not a superposition), since components with different $m$ are generated by the jump term in the master equation, not by the unitary part. Using the mean-field states in Eq. (50) with different $\varphi$, we can construct such a statistical mixture as
$\hat{\rho}_{N}\propto\int d\varphi|\Psi_{\mathrm{MF}}^{(N)}(\varphi)\rangle\langle\Psi_{\mathrm{MF}}^{(N)}(\varphi)|,$ (52)
since the off-diagonal elements with different $m$ and $m^{\prime}$ in the integrand are proportional to $e^{i(m^{\prime}-m)\varphi}$ and vanish by integration. The state in Eq. (52) also has the same distribution as Eq. (51).
The dotted lines in Fig. 1(c) show the distribution in Eq. (51). We find that $Q_{m}$ in the $N=8$ subspace (also in the subspaces with smaller $N$) agrees well with Eq. (51), which indicates that the steady state reached after a long time is well approximated by a direct sum of Eq. (52) with different $N$. This result is consistent with the experimental and mean-field results in Ref. [21], in which fully polarized states in the $x$-$y$ directions were obtained after time evolution.
By contrast, as shown in Fig. 1(c), the distributions in the subspaces with larger particles ($N=18$ and $20$) are narrower than Eq. (51). In particular, for $N=N_{\mathrm{ini}}=20$, the distribution $Q_{m}=\delta_{m,0}$ remains unchanged from the initial state. However, this state $\rho_{N=20}$ obtained as the steady state in the $N=20$ subspace is strikingly different from the initial state in Eq. (42). This difference is easily understood from the simple two-particle example, in which the initial state in Eq. (16) contains various states, whereas the steady state in Eq. (22) only contains the maximum spin state in the $N=N_{\mathrm{ini}}$ subspace. The $N=N_{\mathrm{ini}}=20$ subspace in the steady state should be
$\rho_{N=20}\propto|F=2N,F_{z}=0\rangle\langle F=2N,F_{z}=0|,$ (53)
according to the proof in Sec. IIIB. This state is highly nonclassical, since it can be written as
$|F=2N,F_{z}=0\rangle\propto\int d\varphi|\Psi_{\mathrm{MF}}^{(N)}(\varphi)\rangle,$ (54)
where the terms with $F_{z}=m$ in the integrand are proportional to $e^{im\varphi}$ and the $m\neq 0$ terms vanish by integration. The state in Eq. (54) is a superposition of the different mean-field states and can be regarded as a Schrödinger-cat-like state. However, the probability of observing this state is very small: $P_{N=20}\sim 10^{-8}$ (see (iii) of Fig. 1(b)).
@@
Figure 2: Time evolution for $q/\hbar=204~\mathrm{Hz}$ with and without quenching of $q$. In the case with quenching, $q$ is suddenly changed to 0 at $t=11.0~\mathrm{ms}$. (a) Transverse magnetization $S_{\perp}$ with quenching (solid line), without quenching (dotted line), and without the quadratic Zeeman term (dashed line). (b) Average particle number $\langle\hat{N}\rangle$ (red or dark gray lines) and normalized total magnetization $F^{2}_{\mathrm{norm}}$ (red or light gray lines) with quenching (solid lines) and without quenching (dotted lines). (c) Particle number distribution at $t=1.76~\mathrm{s}$ with quenching (filled boxes) and without quenching (striped boxes).
@@ IV.2 With quadratic Zeeman effect
As shown in the previous subsection, the steady state in the $N=N_{\mathrm{ini}}=20$ subspace is the highly nonclassical state $|F=2N,F_{z}=0\rangle$. In general, deviations from the mean-field approximation become more significant in subspaces with larger $N$. However, the survival probability $P_{N}$ for the steady state is extremely low for large $N$, as shown in Fig. 1(b), and experimental observation is almost impossible.
To enhance $P_{N}$ for the case with large $N$ in the steady state, we include the quadratic Zeeman effect. As has been shown in Ref. [21], the quadratic Zeeman effect changes the relative phases between different magnetic sublevels, which can increase the transverse magnetization. The increase in the transverse magnetization enhances $P_{N}$ for large $N$, since the atomic spins tend to align in the same direction and atomic loss is suppressed.
@@
We include the quadratic Zeeman term with $q/\hbar=204~\mathrm{Hz}$ in the Hamiltonian in Eq. (8), which corresponds to a magnetic field of $484~\mathrm{mG}$. This value of $q$ is chosen so that the initial rise in the transverse magnetization is maximized. The dotted line in Fig. 2(a) shows the transverse magnetization $S_{\perp}$. In the early stage of evolution, $S_{\perp}$ rapidly increases because of the quadratic Zeeman effect, and then oscillates. Such behavior has also been observed experimentally [21]. Subsequently, $S_{\perp}$ decays to a small value because of the particle loss. The dotted lines in Fig. 2(b) show the time evolution of the average particle number $\langle\hat{N}\rangle$ and the normalized total magnetization $F^{2}_{\mathrm{norm}}$. The value of $F^{2}_{\mathrm{norm}}$ first oscillates and then increases toward unity. Although the final steady state should have $F^{2}_{\mathrm{norm}}=1$ according to the proof in Sec. IIIB, the convergence is slow. The average particle number $\langle\hat{N}\rangle$ monotonically decreases with time and converges to a small value, since the steady state consists of only a limited set of states, as given by Eqs. (40) and (41). Consequently, the probability $P_{N}$ is localized near $N=2$, and $P_{N}$ for large $N$ is still very small in the steady state [Fig. 2(c)]. In this scheme of a constant $q$, therefore, the subspaces with $N\simeq N_{\mathrm{ini}}$ do not survive and a nonclassical state cannot be obtained.
To obtain the nonclassical state, we quench the quadratic Zeeman coefficient to $q=0$ at $t=11.0~\mathrm{ms}$, at which $S_{\perp}$ reaches its first peak. Quenching of $q$ is realized by switching off the external magnetic field. The solid lines in Figs. 2(a) and 2(b) show that $\langle\hat{N}\rangle$ and $S_{\perp}$ remain large for a long time as a result of quenching of $q$. This is attributed to the survivable states being extended from Eqs. (40) and (41) to Eq. (24). Figure 2(b) shows that $F^{2}_{\mathrm{norm}}$ converges to unity, which indicates that the fully magnetized steady state is reached. Figure 2(c) shows that, in the steady state, $P_{N}$ has a peak at $N=18$ and the probability $P_{N_{\mathrm{ini}}}=P_{20}$ that no atoms are lost is $\simeq 0.13$, which is much larger than that in the case without quenching of $q$. Thus, the nonclassical state in Eq. (53) is obtained with a probability of $13\%$ by this quench protocol. As shown in Fig. 2(c), the steady state consists not only of the $N=20$ subspace but also of subspaces with other $N$. If we can detect the lost atoms escaping from the trap, the particle number $N$ remaining in the system can be determined and we obtain the pure nonclassical state in Eq. (54).
@@ V Conclusions
We have investigated the dynamics of spin-2 bosons with atomic dissipation. In a previous related study [21], it was shown experimentally and numerically that dissipation can enhance the transverse magnetization in the time evolution, and the results of the mean-field calculations were found to agree well with the experiments. By contrast, the present work focused on the quantum many-body nature of a system with a small number of particles. We proved that the steady states for this system are restricted to statistical mixtures of states with maximum total spins, given by Eq. (26) with Eqs. (24) and (25). We performed numerical simulations using the master equation with and without the quadratic Zeeman term. In the simulations without the quadratic Zeeman term, we confirmed that the system indeed reached a steady state with the maximum total spins. However, the probability of obtaining the nonclassical state in Eq. (53) was extremely small. In the simulations that included the quadratic Zeeman term, the transverse magnetization was enhanced, but the nonclassical state was still lost from the final steady state. When the quadratic Zeeman term was removed in the early stages of time evolution, a relatively large number of particles was found to remain in the steady state. Therefore, the nonclassical state in Eq. (53) can be obtained with a high probability. This finding suggests that particle dissipation not only destroys the quantum properties but also highlights nontrivial quantum many-body states. Although we considered $^{87}\mathrm{Rb}$ atoms in the present study, other atomic species, such as $^{23}\mathrm{Na}$ atoms, may be experimentally suitable, since their inelastic decay rates are much larger [28, 29, 30] and steady states can be reached much faster.
Acknowledgements.
@@
Appendix A Distribution $Q_{m}$ of $z$-magnetization in subspaces with different NN
Figure A.1: Distribution $Q_{m}$ of $z$-magnetization in subspaces with different particle numbers $N$ at final stage of evolution. The dotted lines show Eq. (51). (a) Distributions $Q_{m}$ for $q=0$ at $t=8.8~\mathrm{s}$, corresponding to Fig. 1 (panels for $N=8$, 18, and 20 are the same as those in Fig. 1(c)). (b) Distributions $Q_{m}$ at $t=1.76~\mathrm{s}$ with quenching from $q/\hbar=204~\mathrm{Hz}$ to 0 at $t=11.0\mathrm{ms}$, corresponding to Fig. 2.
In this Appendix, we present the $z$-magnetization distribution $Q_{m}$ in each subspace for different particle numbers $N$ at the final stage of evolution. In Fig. A.1, the histograms represent the distributions $Q_{m}$ obtained from the numerical simulations, whereas the dotted lines plot Eq. (51). Figure A.1(a) shows the results for $q=0$ with the initial state in Eq. (42). The distributions agree well with Eq. (51) for small NN ($6\leq N\leq 10$). By contrast, for large $N$, the distributions tend to localize around $m=0$. However, the probability of occupying these subspaces is very small, as discussed in the main text. Figure A.1(b) shows the distributions after quenching from $q/\hbar=204~\mathrm{Hz}$ to $q=0$. The distributions for small NN ($4\leq N\leq 8$) are well described by Eq. (51). As in the $q=0$ case, the distributions for large $N$ are localized around $m=0$; however, they appear with substantially higher probabilities than those for $q=0$ (see Figs. 1(b) and 2(b)). In the $N=N_{\rm ini}=20$ subspace, the states are strictly restricted to the $F_{z}=0$ sector. These states are described by Eqs. (53) and (54).